OpenStudy (anonymous):

Solve quadratic equation by factoring (without using quadratic formula) and give exact values as answers.

OpenStudy (anonymous):

\[6a^2-17a+12=0\]

OpenStudy (anonymous):

Please show individual steps as to how you would factor that without using quadratic formula.

OpenStudy (australopithecus):

https://www.youtube.com/watch?v=EeVqtpuMFOU

OpenStudy (australopithecus):

I highly recommend watching this video and his others for pre-calculus help

OpenStudy (anonymous):

WITHOUT USING QUADRATIC FORMULA. The video is useless. It shows quadratic formula, I already know it. I am asking how you would do it without using quadratic formula @Australopithecus

OpenStudy (australopithecus):

use completing the square method then

OpenStudy (australopithecus):

https://www.youtube.com/watch?v=xGOQYTo9AKY

OpenStudy (australopithecus):

or you can just factor

OpenStudy (australopithecus):

I would use completing the square method

OpenStudy (anonymous):

How can you just factor? That's what I am asking. How would you do that.

OpenStudy (shayaan_mustafa):

Hi genius12 :) How are you dear?

OpenStudy (shayaan_mustafa):

\[\Large 6a^2 -9a-8a+12=0\]

OpenStudy (ujjwal):

=6a^2-8a-9a+12 =2a(3a-4)-3(3a-4) =(3a-4)(2a-3)